ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 153, 2006

Adler, J. D. and D. Prasad On certain multiplicity one theorems  221
Ben-Yaacov, I. Schrödinger’s cat  157

Blekherman, G. There are significantly more nonnegative polynomials than sums of squares

355

Dahmani, F. Accidental parabolics and relatively hyperbolic groups  93

Darmon, H. and R. Pollack Efficient calculation of Stark–Heegner points via overconvergent modular symbols

319

Fernando, J. F. and J. M. Gamboa Polynomial and regular images

of ℝ^{n} 61

Gamboa, J. M. See J. F. Fernando

Huttner, M. Constructible sets of linear differential equations and effective rational approximations of polylogarithmic functions

1

Ito, S. and H. Rao Atomic surfaces, tilings and coincidence I. Irreducible case

129

Lubinsky, D. S. Jackson and Bernstein theorems for the weight

exp(−|x|) on ℝ 193

Mankiewicz, P. and N. Tomczak-Jaegermann Low dimensional sections versus projections of convex bodies

45

Montenegro, R. A sharp isoperimetric bound for convex bodies  267
Parker, C. and J. Saxl Localization and finite simple groups  285
Parker, C. and J. Saxl Two maximal subgroups of E_{8}(2)  307

Pollack, R. See H. Darmon

Prasad, D. See J. D. Adler

Rao, H. See S. Ito

Ryazanov, V., U. Srebro and E. Yakubov Finite mean oscillation and the Beltrami equation

247

Saxl, J. See C. Parker

Srebro, U. See V. Ryazanov

Tomczak-Jaegermann, N. See P. Mankiewicz

Yakubov, E. See V. Ryazanov